# EASSE-DE: Easier Automatic Sentence Simplification Evaluation for German

### v1

@@ 4.1.1 German TS Test Sets
[…]
@@ 4.1.2 German TS Models
For German TS, a few models are available or reproducible, e.g., ZEST, by Mallinson et al. (2020), sockeye by Spring et al. (2021), custom-decoder-ats by Anschütz et al. (2023), the few-shot approaches on BLOOM by Ryan et al. (2023), or the mBART models by Stodden et al. (2023). A more detailed description and analysis of German TS models, including their reproduction, has been recently proposed by Stodden (2024). The system outputs of all reproduced German TS models (see Table 2) have been added to EASSE-DE Table 2 to facilitate a better comparison between existing models and models which will be newly proposed in future.
@@
hda-etr Siegel et al. (2019) rule-based - - https://github.com/hdaSprachtechnologie/easy-to-understand_language
@@
sockeye-DEplain-APA - seq2seq DEplain-APA 10,660 https://huggingface.co/DEplain
@@
DEplain-APA 10,660 https://huggingface.co/DEplain/trimmed_mbart_sents_apa
@@
DEplain-APA 10,660 https://huggingface.co/DEplain
@@
SGC 4,430 https://huggingface.co/DEplain
@@
- - https://github.com/XenonMolecule/MultiSim
@@
200 & 959 https://github.com/XenonMolecule/MultiSim
@@
200 & 959 https://github.com/XenonMolecule/MultiSim
@@
Table 2: Overview of German TS models including training details (i.e., training data and size of training samples). Each line separates different model types. Adaptation from Stodden (2024).
@@ 4.2.1 Method
Further, we have selected the generated simplifications of one model, i.e., mBART-DEplain-APA+web. Reasons for the choice of this model are that it is ready-to-use without additional examples, and, following Stodden (2024), this model achieves the best BERT-Scores across several test sets. In comparison, the BLOOM models by Ryan et al. (2023) are few-shot models that require additional complex-simple pairs to generate simplifications.
@@ 4.2.2 Results
[…]
@@
As expected, different tokenization methods (including language specification) affect the calculation of metrics used for TS evaluation. The last three rows in each block of Table 3 show the differences in the scores when using different tokenization strategies. We can see that the BERT-score is always the same for all settings due to the sub-word tokenization in BERT. The FRE scores are also robust across all test sets when looking at the trimmed-mBART results, but in Appendix A Table 4, we see slightly more differences. The SARI scores also change slightly, i.e., to less than 1 point in all settings, whereas the differences in the BLEU scores range between 2 to 3 points in all test sets. In conclusion, when comparing one model against another with a slightly different evaluation setting (here, the tokenizer), even these small changes can be wrongly interpreted as an improvement of the model idiosyncrasy. However, it is only due to the different settings. Therefore, we recommend stating all settings chosen for evaluation for a more reliable comparison between models.
@@ Appendix A Results of Identity Baseline
[…]

### v2

@@ 4.1.1 German TS Test Sets
name target group domain size # ref. n:m FREnormal-$\downarrow$ sent. len.normal-$\uparrow$ word len.normal-$\uparrow$ FREnormal-$\uparrow$ sent. len.normal-$\downarrow$ word len.normal-$\downarrow$
@@ 4.1.2 German TS Models
For German TS, a few models are available or reproducible, e.g., ZEST, by Mallinson et al. (2020), sockeye by Spring et al. (2021), custom-decoder-ats by Anschütz et al. (2023), the few-shot approaches on BLOOM by Ryan et al. (2023), or the mBART models by Stodden et al. (2023). A more detailed description and analysis of German TS models, including their reproduction, has been recently proposed by Stodden (2024 (to appear). The system outputs of all reproduced German TS models (see Table 2) have been added to EASSE-DE Table 2 to facilitate a better comparison between existing models and models which will be newly proposed in future.
@@
hda-etr Siegel et al. (2019) rule-based - -
https://github.com/hdaSprachtechnologie/easy-to-understand_language
@@
sockeye-DEplain-APA - seq2seq DEplain-APA 10,660
https://huggingface.co/DEplain
@@
DEplain-APA 10,660
https://huggingface.co/DEplain/trimmed_mbart_sents_apa
@@
DEplain-APA 10,660
https://huggingface.co/DEplain
@@
SGC 4,430
https://huggingface.co/DEplain
@@
https://github.com/XenonMolecule/MultiSim
@@
200 & 959
https://github.com/XenonMolecule/MultiSim
@@
200 & 959
https://github.com/XenonMolecule/MultiSim
@@
Table 2: Overview of German TS models including training details (i.e., training data and size of training samples). Each line separates different model types. Adaptation from Stodden (2024 (to appear).
@@ 4.2.1 Method
Further, we have selected the generated simplifications of one model, i.e., mBART-DEplain-APA+web. Reasons for the choice of this model are that it is ready-to-use without additional examples, and, following Stodden (2024 (to appear), this model achieves the best BERT-Scores across several test sets. In comparison, the BLOOM models by Ryan et al. (2023) are few-shot models that require additional complex-simple pairs to generate simplifications.
@@ 4.2.2 Results
Tok. Lang. BLEUnormal-$\uparrow$ SARInormal-$\uparrow$ BS-Pnormal-$\uparrow$ FREnormal-$\uparrow$
@@
As expected, different tokenization methods (including language specification) affect the calculation of metrics used for TS evaluation. The last three rows in each block of Table 3 show the differences in the scores when using different tokenization strategies. We can see that the BERT-score is always the same for all settings due to the sub-word tokenization in BERT.
The FRE scores are also robust across all test sets when looking at the trimmed-mBART results, but in Appendix A Table 4, we see slightly more differences. The SARI scores also change slightly, i.e., to less than 1 point in all settings, whereas the differences in the BLEU scores range between 2 to 3 points in all test sets. In conclusion, when comparing one model against another with a slightly different evaluation setting (here, the tokenizer), even these small changes can be wrongly interpreted as an improvement of the model idiosyncrasy. However, it is only due to the different settings. Therefore, we recommend stating all settings chosen for evaluation for a more reliable comparison between models.
@@ Appendix A Results of Identity Baseline
Tok. Lang. BLEUnormal-$\uparrow$ SARInormal-$\uparrow$ BS-Pnormal-$\uparrow$ FREnormal-$\uparrow$